A compound thin cylinder has a common diameter of 100 mm and the inner cylinder has a thickness of 2,5 mm. The radial pressure between the two cylinders is 200 kPa and the difference between the two common diameters before shrinkage was 4,305 x 10-3 mm. Determine (a) the thickness of the outer cylinder; and (b) the resultant hoop stresses in both cylinders if the compound cylinder is subjected to an internal pressure of I80 kPa (E = 200 GPa). 


1
Expert's answer
2021-06-18T07:55:05-0400

Part a

"\sigma _h= \frac{Pd}{2t}"

"\in_n=\frac{\sigma_n}{\in}-\frac{Pd}{2 \in \sigma}"

"\frac{\triangle d}{d}=\in_h-\frac{Pd}{2t \in} \implies \triangle d= \frac{Pd^2}{2t \in}"

"\frac{Pd_1^2}{2t_1 \in} +\frac{Pd_0^2}{2t_0 \in} =\triangle d_1+\triangle d_0"

"\frac{P}{2 \in}[\frac{100^2}{\in_1}+\frac{95^2}{2.5}]=\frac{4.305}{10^2}"

"[\frac{100^2}{t_1}+3610]=\frac{4.305}{10^3}* \frac{2*200*10^3}{0.02}"

"t_1=\frac{100^2}{5000}=2 mm"

Thickness of the outer cylinder is 2 mm


Part b

Consdering the inner cylinder

"\sigma_n=\frac{Pd}{2t}=\frac{(180-200)*10^3*95}{2*2.5}=-380 kPa"

Therefore "\sigma_n = 380 kPa" (Compresive)